# The Nontrivial Vacuum Structure of an Extended $t\overline{t}$ BEH (Higgs) Bound state

### v1

@@ Abstract
In a recent reformulation of top-quark condensation for the Brout-Englert-Higgs boson, we introduced an extended internal wave-function, $\phi(r)$. We show how this leads to a manifestly Lorentz invariant formalism, where the absence of “relative time” is a gauge invariance of the bilocal field theory. This dictates a novel and nontrivial Lorentz invariant vacuum structure for the BEH boson, the relativistic generalization of a condensed matter state such as a BCS or Bose-Einstein condensate.
@@ I Introduction
In the case of a Brout-Englert-Higgs (BEH) boson the constituents are massless chiral top and anti-top quarks located at space-time coordinates $(x^{\mu},y^{\mu})=X^{\mu}\pm r^{\mu}$. $H(X^{\mu})$ can be viewed as the standard model (SM) BEH isodoublet with electroweak charges and $\phi(r^{\mu})$ is a complex scalar which is electroweak neutral chris1 ; chris2 . $H(X^{\mu})$ describes the center-of mass motion of the BEH boson, and $\phi(r^{\mu})$ is then the internal wave-function.
@@
A theory of a BEH boson composed of $\overline{t}t$, known as “top condensation,” was proposed in the 1990’s Yama ; BHL ; Topcolor ; NSD , deploying the Nambu-Jona-Lasinio (NJL) model NJL with renormalization group (RG) improvements BHL . The NJL model, however, is pointlike, lacking $\phi(r^{\mu})$, which leads to difficulties when there is a large hierarchy between the composite scale, $M_{0}$, and the electroweak scale $|\mu|$ (the symmetric phase mass of the BEH boson, $|\mu|=88$ GeV). The inclusion of the internal wave-function yields significant improvement in the predictions of the low energy parameters. In the limit $|\mu|<\!\!<M_{0}$ there is significant wave-function spreading of $\phi(\vec{r})$ and the resulting dilution effects dominate the low energy effective theory. This brings its predictions into concordance with experiment, virtually eliminating fine-tuning, and predicting the new mass scale of the binding interaction, $M_{0}\sim 6$ TeV, ref.(chris1 ,chris2 ) summarized in Appendix A.
@@
We emphasize that the theory is manifestly Lorentz invariant. The challenge is, however, that the internal wave-function, $\phi(r^{\mu})$, introduces a priori unwanted dependence upon “relative time.” This is $r^{0}=(x^{0}-y^{0})/2$ in the rest frame of the bound state, but boosted relative time would occur in any frame. Due to the single time parameter of Hamiltonian based quantum mechanics, the relative time is unphysical and must be removed in a manner consistent with Lorentz invariance. This issue is avoided in the NJL model due to the pointlike interaction, but the relative time problem is well known to arise in any bound state with an extended interaction, Dirac . It can be seen as arising in the free field limit by kinematics (below).
@@
Relative time implies an “arrow of relative time,” a timelike, unit, 4-vector $\omega_{\mu}$, associated with $\phi(r^{\mu})$. The relative time is then $\tau$, where $r^{\mu}=\omega^{\mu}\tau$. The absence of relative time can then be viewed as a gauge symmetry of the internal wave-function $\phi(r^{\mu})$, where a “gauge transformation” is $r^{\mu}\rightarrow r^{\mu}+\omega^{\mu}\tau$ with $\tau$ acting as a gauge parameter. Given an $\omega^{\mu}$ we can pass to a manifestly “gauge invariant” $\phi(r^{\mu})$ field, the analogue of a“Stueckelberg” field, where the symmetry is built in:
$\phi(r^{\mu})\rightarrow\phi_{\omega}(r^{\mu})\equiv\phi\left(\omega^{\mu}(\omega_{\nu}r^{\nu})-r^{\mu}\right)\qquad\omega^{2}\equiv\omega_{\mu}\omega^{\mu}=1.$ (2)
The issue then becomes: “what defines $\omega_{\mu}$?”
@@
For simple two-body bound states, where $H(X)\sim\exp(iP_{\mu}X^{\mu})$, then $\omega^{\mu}$ can be identified with the normalized 4-momentum, $\omega^{\mu}=P^{\mu}/\sqrt{|P^{2}|}$ (e.g. via Lagrange multipliers). Hence, a two-body spherically symmetrical bound state becomes:
$H(X)\phi_{\omega}(r^{\mu})=\exp(iP_{\mu}X^{\mu})\;\phi\!\left(\sqrt{((P_{\mu}r^{\mu})^{2}/P^{2})-r_{\mu}r^{\mu}}\right)$ (3)
@@
$\phi(r^{\mu})$ then reverts to $\phi(\vec{r})$ in the rest frame where it can be treated as a solution to a static Schrödinger-Klein-Gordon (SKG) equation chris1 .
@@
However, there remains the challenging question when $\mu^{2}<0$ and spontaneous symmetry breaking occurs: “How can we have a Lorentz invariant vacuum state with $P^{\mu}=0$, but nonzero $\omega_{\mu}$?” That is, “what determines $\omega^{\mu}$ in the vacuum?” Any constraint that locks $P^{\mu}$ to $\omega^{\mu}$ in the two-body case ceases to exist in the vacuum and $\omega^{\mu}$ would seem to become arbitrary. While one possibility is that $\phi(r^{\mu})\rightarrow$(constant) in the vacuum, this does not lead to a consistent result with the internal dynamics of $\phi(r)$.
@@
It would seem nonsensical to assert that the vacuum is defined by a condensate that occurred in a particular Lorentz frame with a particular $\omega_{\mu}$. If, for example, $\omega_{\mu}$ is somehow associated with the local cosmic rest frame we would obtain induced Lorentz violating effects in the electroweak physics, e.g., vacuum Cerenkov radiation for all particles that receive mass from the BEH boson, top-quarks to electrons and neutrinos ColemanGlashow . We would also have potentially large radiatively induced Lorentz non-invariant corrections to electrodynamics Kostelecky (we have estimated these effects in chris1 and while they are suppressed as $\sim|\mu|^{2}/M_{0}^{2}$, they appear nonetheless to be problematic). We would prefer a starting point in which the vacuum is manifestly Lorentz invariant and contains no preferred $\omega^{\mu}$. Nonetheless, we require $\omega_{\mu}$ to define the solutions, $\phi_{\omega}(r^{\mu})$, to implement the relative time invariance of our theory.
@@
Therefore, we propose the following solution to the vacuum problem: the vacuum is a Lorentz invariant sum over all frames of the individual solutions $\phi_{\omega}(r^{\mu})$ in each frame. We will introduce a Lorentz invariant integral over $\omega^{\mu}$, leading to a novel internal wave-function for the vacuum state, $\Phi(r^{\mu})$, with (unrenormalized) $H^{\prime}(X)$.
$H(x,y)=H^{\prime}(X)\Phi(r)\qquad\text{where,}\qquad\Phi(r)={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})$ (4)
@@
With this definition of the vacuum, $\Phi(r)$ is a collective state, similar to a condensed matter system, such as a BCS superconductor, BCS , or Bose-Einstein condensate (BEC) in, e.g., the Gross-Pitaevskii model, or Ginzburg-Landau model of superconductivity BEC . However, our new vacuum differs from these because it is Lorentz invariant.
For example, in a BEC, created in a laboratory, one coaxes a large number, $N$, of bosons into a state with common 3-momentum, such as $\vec{P}=0$. Hence two-body scattering states are defined by the relative momenta, $\small\vec{Q}_{i}$, and would have bilocal wave-functions, $\small\sim a_{i}e^{i\vec{Q}_{i}\cdot\vec{r}}$. The Hamiltonian kinetic term for the pairs is then $\small\sim\sum_{i}|a_{i}|^{2}Q^{2}_{i}\propto N$ while the short-distance, two-body, attractive scattering potential is $\small\sim-M\sum_{i}\sum_{j}a^{\dagger}_{i}a_{j}\propto N^{2}$. In the large $N$ limit the Hamiltonian is then dominated by the two-body scattering effects and is diagonalized by a coherent state that is a sum over all $\vec{Q}_{i}$ as , $\small\Phi\sim\sum_{i}a_{i}e^{i\vec{Q}_{i}\cdot\vec{r}}/\sqrt{N}$. This dominance of the interaction is a key feature of condensates. It occurs in the BCS state of a superconductor (analogously the Ginzburg-Landau wave-function), that is a sum (or product in kets) over all Cooper pairs with $\vec{Q}_{i}$ lying on the Fermi surface in momentum space, BEC ; BCS . These states exist in a preferred, $\vec{P}=0$ frame, the rest frame of the system.
@@
The vacuum correlation function we propose presently, $\Phi(r)$, is a sum over constituent wave-function solutions of the SKG equation, each with $P^{\mu}=0$ but nontrivial $\phi_{\omega}(r^{\mu})$, each specified by a different Lorentz frame, $\omega_{\mu}$. The $\omega_{\mu}$ are timelike and span the timelike hyperboloid in momentum space. The $\phi_{\omega}(r^{\mu})$, in a sense, are the relativistic generalization of “Cooper pairs” BCS . While Cooper pairs span the Fermi surface in a superconductor, the $\Phi\sim\int_{\omega}\phi_{\omega}$ span the future timelike hyperboloid in the 4-vectors, $\omega$. It follows that the invariant integral over $\omega_{\mu}$ yields a Lorentz invariant $\Phi(r)$.
@@
Figure 1: Vacuum wave-function, $\Phi(r)$, spanning the timelike hyperboloid in 4-vectors $\omega_{\mu}$ by integrating over “internal wave-functions,” $\phi_{\omega}(r)$, to form the Lorentz invariant $\Phi(r)={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})$.
Though the component fields, $\phi_{\omega}(r)$, satisfy a nontrivial integro-differential equation with dependence upon $\omega^{\mu}$, there is no requirement on the relative time in the vacuum (any Lagrange multipliers that lock $\omega^{\mu}$ to a $P^{\mu}$ are now absent). Upon integrating over $\omega^{\mu}$ the time component of $r^{\mu}$ returns in $\Phi(r^{\mu})$, but the arrow of relative time, $\omega^{\mu}$, is removed. The BEH boson observed at the LHC, $h(X,r)$ and the Nambu–Goldstone phases that become longitudinal $W^{\pm}$ and $Z^{0}$ emerge as an “excitons” of this collective state, e.g., as $h(X,r)\propto h(X)\Phi(r^{\mu})$. While dependence upon $r^{0}$ is now a feature of the vacuum wave-function, $\Phi(r)$, analogous to a vacuum two-point function, or a propagator, this makes no reference to a particular $\omega_{\mu}$.
The presence of $\Phi(r)$ is unobservable in the kinetic terms of the BEH boson, gauge bosons and top quarks, which conform exactly to the SM. We do, however, have access to its structure in the Yukawa and quartic interactions through an expansion in powers of $r^{\mu}$, which corresponds to an expansion in inverse powers of $M_{0}$. This leads to a series of higher dimension operators, generating novel, Lorentz invariant, effects that are small. In the present paper we will only briefly describe how these effects can be extracted from the Yukawa coupling in the Section IV (see also chris1 ). We presently ignore the complications of the introduction of all flavors of quarks and leptons, in particular the $b_{R}$ quark. We expect these to be perturbative and follow the earlier papers on extended technicolor Technicolor and “topcolor” from the 1990’s Topcolor . In this paper we will mainly discuss the nontrivial vacuum.
@@
This theory of the vacuum emerges from the underlying theory of two-body bound states with $H^{\prime}(X)\Phi(r)=H(X)\widetilde{\Phi}(r)$ where $\widetilde{\Phi}(r)\propto\Phi(r)$ is the classical average over all $\phi_{\omega}(r^{\mu})$ and $\langle H\rangle=v_{weak}$. In particular we have,
$\widetilde{\Phi}(0)=\phi_{\omega}(0)\qquad\text{for any $\omega_{\mu}$.}$ (5)
This relationship establishes that the predictions of the full theory $M_{0}$, $\lambda$, etc., with the full vacuum structure $\Phi(r)$ included, agree with the earlier results from the simple two-body bound state studies in chris1 , and are valid to order $|\mu|^{2}/M_{0}^{2}$. In the broken phase we then have:
@@
$H^{\prime}(X)\Phi(r)=H(X)\widetilde{\Phi}(r^{\mu})\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}+\frac{1}{\sqrt{2}}h(X)\\ 0\end{array}\biggr)\widetilde{\Phi}(r^{\mu})$ (8)
where $h(X)$ is the “Higgs field” of the standard model. The $H^{\prime}(X)\Phi(r)$ action will involve renormalized parameters but will be seen to be completely consistent with the two-body $H(X)\widetilde{\Phi}$ action, classically averaging over $\phi_{\omega}$. Here the constant zero 4-momentum VEV, $v_{weak}$, is carried by $H(X)$ and determined in the usual way by the minimum of the sombrero potential.
@@
The effective action of the $H^{\prime}(X)\Phi(r)$ bilocal field is completely consistent with the simpler two-body $H(X)\phi_{\omega}(r)$ action. The main prediction is the existence of the new binding interaction at the scale $M_{0}\approx 6$ TeV and the emergence of an octet of colorons coupled most strongly to third generation quarks. This interaction is only partially strong (approximately half critical) since the critical behavior occurs only in the binding channel where loop effects reinforce the binding interaction chris2 .
II Absence of Relative Time as a Gauge Invariance

### v3

@@ Abstract
In a recent reformulation of top-quark condensation for the Brout-Englert-Higgs boson, we introduced an extended internal wave-function, $\phi(r)$. We show how this leads to a manifestly Lorentz invariant formalism, where the absence of “relative time” is a gauge invariance of the bilocal field theory. This dictates a novel and nontrivial Lorentz invariant vacuum structure for the BEH boson, the relativistic generalization of a condensed matter state analogous to a BCS condensate.
@@ I Introduction
In the case of a Brout-Englert-Higgs (BEH) boson, composed of $\overline{t}t$, massless chiral top and anti-top quarks are located at space-time coordinates $(x^{\mu},y^{\mu})=X^{\mu}\pm r^{\mu}$. $H(X^{\mu})$ can be viewed as the standard model (SM) BEH isodoublet with electroweak charges and $\phi(r^{\mu})$ is a complex scalar which is electroweak neutral chris1 ; chris2 . $H(X^{\mu})$ describes the center-of mass motion of the BEH boson, and $\phi(r^{\mu})$ is then the internal wave-function.
@@
A theory of a BEH boson composed of $\overline{t}t$, known as “top condensation,” was proposed in the 1990’s Yama ; BHL ; Topcolor ; NSD , deploying the Nambu-Jona-Lasinio (NJL) model NJL with renormalization group (RG) improvements BHL . The NJL model, however, is pointlike, lacking $\phi(r^{\mu})$, which leads to difficulties when there is a large hierarchy between the composite scale, $M_{0}$, and the electroweak scale $|\mu|$ (the symmetric phase mass of the BEH boson, $|\mu|=88$ GeV). The inclusion of the internal wave-function yields significant improvement in the predictions of the low energy parameters. In the limit $|\mu|<\!\!<M_{0}$ there is significant wave-function spreading of $\phi(\vec{r})$ and the resulting dilution effects dominate the low energy effective theory. This brings its predictions into concordance with experiment, virtually eliminating fine-tuning, and predicting the new mass scale of the binding interaction, $M_{0}\sim 6$ TeV, ref.(chris1 ,chris2 ) summarized in Appendix B.
@@
We emphasize that the theory is manifestly Lorentz invariant. The challenge is, however, that the low energy internal wave-function, $\phi(r^{\mu})$, introduces a priori unwanted dependence upon “relative time.” This is $r^{0}=(x^{0}-y^{0})/2$ in the rest frame of the bound state, but boosted relative time would occur in any frame. Due to the single time parameter of Hamiltonian based quantum mechanics, the relative time is unphysical in a wave-function.
The wave-function can be viewed as the “end-cap” of the path integral that begins (and terminates) on given time-slices. The path-integral is a Green’s function of the Hamiltonian–Schrödinger equation that propagates the wave-function from an initial time and spatial configuration $(t_{0},\vec{X}_{0},\vec{r})$ to a future $(t_{1},\vec{X}_{1},\vec{r}_{1})$ (this is true in field theory where the initial and final configurations are static fields, $\phi(\vec{x}_{i})$, specified on given time slices, $t$). The relative time does not exist on these end-cap time slices and plays no role in the initial data. [^1] We remind the reader that, for the Hydrogen atom with potential, $V(\vec{r})$ and Schrödinger equation $H\psi(\vec{r},t)=i\partial_{t}\psi(\vec{r},t)$, the parameters $t$ and $\vec{r}$ do not form a 4-vector under Lorentz covariance. Rather, $(t,\vec{X})\rightarrow X^{\mu}$ becomes a relativistic 4-vector, with $\vec{X}$ describing the center-of-mass of the atom, and $\vec{r}\rightarrow r^{\mu}$ with the relative time $r^{0}$ constrained to zero. Hence, for wave-functions, relative time must be removed in a manner consistent with Lorentz invariance. The issue of relative time is avoided in the NJL model due to the pointlike interaction, but the absence of relative time is a well known challenge known to arise in any bound state with an extended interaction, Dirac .
@@
The problem of maintaining Lorentz invariance is that relative time implies an “arrow of relative time,” a timelike, unit, 4-vector $\omega^{\mu}$, associated with $\phi(r^{\mu})$. The relative time is then $\tau$, where $r^{\mu}=\omega^{\mu}\tau$. The absence of relative time can then be viewed as a gauge symmetry of the internal wave-function $\phi(r^{\mu})$, where a gauge transformation is $r^{\mu}\rightarrow r^{\mu}+\omega^{\mu}\tau$ with $\tau$ acting as a gauge parameter. Given an $\omega^{\mu}$ we can then pass to a manifestly gauge invariant $\phi_{\omega}(r^{\mu})$ field, the analogue of a “Stueckelberg” field, where the symmetry is built in:
$\phi(r^{\mu})\rightarrow\phi_{\omega}(r^{\mu})\equiv\phi\left(\omega^{\mu}\omega_{\nu}r^{\nu}-r^{\mu}\right)\qquad\omega^{2}\equiv\omega_{\mu}\omega^{\mu}=1.$ (2)
@@
The issue then becomes how to avoid Lorentz non-invariance with an arbitrary input $\omega^{\mu}$, in particular, “what defines $\omega^{\mu}$?” For simple two-body bound states, where $H(X^{\mu})\sim\exp(iP_{\mu}X^{\mu})$, then $\omega^{\mu}$ can be identified with the normalized 4-momentum, $\omega^{\mu}=P^{\mu}/\sqrt{P^{2}}$ where $P^{2}=P_{\mu}P^{\mu}=\mu^{2}>0$. Hence, a two-body spherically symmetrical bound state becomes:
$H(X^{\mu})\phi_{\omega}(r^{\mu})=\exp(iP_{\mu}X^{\mu})\;\phi\!\left(\sqrt{((P_{\mu}r^{\mu})^{2}/P^{2})-r_{\mu}r^{\mu}}\right).$ (3)
@@
This state is then explicitly Lorentz invariant. Then $\phi(r^{\mu})$ reverts to $\phi(|\vec{r}|)$ in the rest frame where it can be treated as a solution to a static Schrödinger-Klein-Gordon (SKG) equation chris1 .
@@
However, there remains the question when $\mu^{2}<0$ and spontaneous symmetry breaking occurs: “How can we have a Lorentz invariant vacuum state with $P^{\mu}=0$, but nonzero $\omega^{\mu}$?” That is, “what determines $\omega^{\mu}$ in the vacuum?” Any constraint that locks $P^{\mu}$ to $\omega^{\mu}$ in the two-body case ceases to exist in the vacuum and $\omega^{\mu}$ would seem to become arbitrary. While one possibility is that $\phi(r^{\mu})\rightarrow$(constant) in the vacuum, this does not lead to a consistent solution to the SKG equation.
@@
It would seem nonsensical to assert that the vacuum is defined by a condensate that occurred in a particular Lorentz frame with a random $\omega^{\mu}$. If, for example, $\omega^{\mu}$ is somehow associated with the local cosmic rest frame we would obtain induced Lorentz violating effects in the electroweak physics, e.g., vacuum Cerenkov radiation for all particles that receive mass from the BEH boson, top-quarks to electrons and neutrinos ColemanGlashow . We have estimated these effects in chris1 and they are suppressed as $\sim|\mu|^{2}/M_{0}^{2}$. While the limit on vacuum Cerenkov radiation for the electron is satisfied, these effects may be problematic, e.g., potentially large radiatively induced Lorentz non-invariant corrections to electrodynamics may arise at loop level Kostelecky . We therefore require a starting point in which the vacuum is manifestly Lorentz invariant, hence it must contain no preferred $\omega^{\mu}$. Yet, we evidently require nonzero $\omega^{\mu}$ to define the solutions, $\phi_{\omega}(r^{\mu})$, to implement the relative time invariance.
@@
Hence, we propose the following solution to the vacuum problem: the vacuum is a Lorentz invariant sum over all frames of the individual solutions $\phi_{\omega}(r^{\mu})$ in each frame. We introduce a Lorentz invariant integral over $\omega^{\mu}$, leading to a novel internal wave-function for the vacuum state, $\Phi(r^{\mu})$, with (unrenormalized) $H^{\prime}(X)$.
$H(x,y)=H^{\prime}(X)\Phi(r^{\mu})\qquad\text{where,}\qquad\Phi(r^{\mu})={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu}).$ (4)
@@
With this definition of the vacuum, $\Phi(r^{\mu})$ becomes a collective state, similar to a condensed matter system such as the BCS superconductor.
Our present calculation is formally similar to the construction of the coherent condensate of Cooper pairs in a BCS superconductor Cooper BCS , differing essentially by our required Lorentz invariance. Cooper pairs are two-body bound states, (approximate) bilocal plane waves in momentum space: $\phi_{i}(\vec{r})\approx\phi_{i}\exp(\vec{k}_{i}\cdot(\vec{x}-\vec{y}))$, where the $\vec{k}_{i}$ lie in a small range immediately above the Fermi surface of the material, hence the “pairing” is an electron of momentum $\vec{k}_{i}$ with the antipodal electron $\vec{k}_{i}$. The Hamiltonian kinetic term for the pairs is then $\small\sim(1/m)\sum_{i}|\phi_{k_{i}}|^{2}k^{2}_{i}\propto N$, while the two-body attractive scattering potential (weak phonon exchange interaction) is $\small\sim-\epsilon m\sum_{i}\sum_{j}\phi^{\dagger}_{k_{i}}\phi_{k_{j}}\propto N^{2}$. Hence the weak phonon interaction is significantly enhanced by large $N\epsilon\sim 1$ and can compete with the kinetic terms to form a stable collective state which forms the condensate. The semiclassical wave-function of the ground state is then $\Phi\sim\sum_{i}^{N}\phi_{i}(\vec{r})$ where the sum spans the Fermi surface. The BCS theory is essentially a bilocal field theory. [^2] The BCS ground state is a coherent state of Cooper pairs, and specified as a product of Dirac kets that are mixtures of vacuum and pairs, $|BCS\rangle\sim\prod^{N_{F}}_{i}(u_{i}|0\rangle+v_{i}|\!\uparrow\!\vec{k}_{i},\!\downarrow\!-\vec{k}_{i}\rangle)$, where the product extends over the $N_{F}$ momenta, $k_{i}$, slightly above the Fermi surface (bounded above by the Debye frequency $\omega_{D}$). The superconductor’s properties are determined by the resulting ratio $v_{i}/u_{i}$, obtained by minimization of the Hamiltonian with, e.g., temperature effects or current dependence, etc. (See e.g., Feynstat Urbana , and related discussion of mean-field approximations GL BEC ). The condensate quantum field is bilocal, with creation operators $a^{\dagger}_{\vec{k}_{i}}$ for electrons of momenta $\vec{k}_{i}$, $\widehat{\Phi}(\vec{r})\sim\sum^{N}_{i}a^{\dagger}_{\vec{k}_{i}}a^{\dagger}_{-{\vec{k}_{i}}}\phi_{i}(\vec{r})$. Note this creates, and sums over, Cooper pairs of antipodal momenta $(\vec{k}_{i},-\vec{k}_{i})$ that span the Fermi surface. The analogue of our semiclassical bilocal field is then $\Phi(\vec{r})=\langle BCS|\widehat{\Phi}(\vec{r})|BCS\rangle\sim\sum^{N}_{i}u_{i}v_{i}^{*}\phi_{i}(\vec{r})$. The ground state (vacuum) expectation value of $\langle\Phi\rangle$ is determined by a gap equation (similar of the original formulation of the NJL model NJL ). A gap equation is equivalent to the minimization of the BEH (Higgs) potential (the “quartic interaction” of the potential is implicit in the gap equation loop, and for us the quartic coupling also arises at loop level).
@@
The Brout-Englert-Higgs (BEH) vacuum wave function which we presently propose, $\Phi(r^{\mu})$, is a sum over all constituent wave-function solutions of the SKG equation, each with $P^{\mu}=0$ but with nontrivial $\phi_{\omega}(r^{\mu})$, each corresponding to a different Lorentz frame, $\omega^{\mu}$, the analogue of the Fermi momenta in the superconductor. The $\phi_{\omega}(r^{\mu})$ are then a relativistic generalization of Cooper pairs BCS , but while Cooper pairs span the Fermi surface in a superconducting condensate, the $\Phi\sim\int_{\omega}\phi_{\omega}$ span the future timelike hyperboloid in the timelike unit 4-vectors, $\omega^{\mu}$.
The component fields, $\phi_{\omega}(r)$, each satisfy a nontrivial integro-differential equation with dependence upon the frame $\omega^{\mu}$. It follows that the invariant integral over $\omega^{\mu}$, normalized by $\cal{N}$}, yields a Lorentz invariant $\Phi(r^{\mu})$. The Hamiltonian is then diagonalized by the coherent state $H(X)\Phi(r^{\mu})$. The normalized effective action for $H(X^{\mu})$ is obtained and, upon integrating out $r^{\mu}$, yielding the “Higgs” potential $\mu^{2}|H|^{2}+(\lambda/2)|H|^{4}$. $H$ then acquires a vacuum expectation value (VEV), determined by a negative SKG eigenvalue, $\mu^{2}$, and stabilized by the loop induced quartic interaction $\lambda$.
@@
Figure 1: Vacuum wave-function, $\Phi(r^{\mu})$, spanning the timelike hyperboloid in 4-vectors $\omega^{\mu}$ by integrating over “internal wave-functions,” $\phi_{\omega}(r)$, to form the Lorentz invariant $\Phi(r^{\mu})={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})$. This is a relativistic analogue of a BCS state that integrates over the Cooper pairs on the Fermi surface.
@@
The vacuum emerges from the underlying theory with $H^{\prime}(X)\Phi(r^{\mu})=H(X^{\mu})\widetilde{\Phi}(r)$ where $\widetilde{\Phi}(r)=(1/\sqrt{N})\Phi(r^{\mu})$ is the classical average over all $\phi_{\omega}(r^{\mu})$ and $\langle H\rangle=v_{weak}$. The BEH boson observed at the LHC, $h(X)$ and the Nambu–Goldstone phases that become longitudinal $W^{\pm}$ and $Z^{0}$, then emerge as “excitons” of the collective state. In the broken phase we then have:
@@
$H(X^{\mu})\widetilde{\Phi}(r^{\mu})\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}+\frac{1}{\sqrt{2}}h(X)\\ 0\end{array}\biggr)\widetilde{\Phi}(r^{\mu}),$ (7)
where $h(X)$ is the “Higgs field” of the standard model.
@@
Here the constant zero 4-momentum VEV, $v_{weak}$, is carried by $H(X^{\mu})$ and determined in the usual way by the minimum of the sombrero potential. The main prediction of the theory remains as the existence of the new binding interaction at the scale $M_{0}\approx 6$ TeV and the emergence of an octet of colorons coupled most strongly to third generation quarks. This interaction is only partially strong (approximately half critical) since the critical behavior occurs only in the binding channel where loop effects reinforce the binding interaction chris2 Jackiw .
In the present paper we will mainly focus on the nontrivial vacuum. We will also briefly sketch how higher dimension operator terms ${\cal{O}}(1/M_{0}^{2})$ may be extracted from the Yukawa coupling in Section IV (see also chris1 ), which may ultimately present observables in high sensitivity flavor physics experiments. We mainly ignore the complications of the introduction of all flavors of quarks and leptons, in particular the $b_{R}$ quark. We expect these to be perturbative and follow the earlier papers on extended technicolor Technicolor and “topcolor” from the 1990’s Topcolor , and will be the subject of future work christodo .
Our main result is a successful, natural, minimally fine-tuned (few $\%$), composite theory of the BEH boson, with a novel physical scale $\sim 6$ TeV corresponding to a new semi-strong interaction within the third generation quarks, and associated gauge fields that are possibly accessible to the LHC (chris1 ,chris2 ). We show that vacuum of this theory maintains the Lorentz invariance, as a dynamical analogue of a BCS superconductor, yielding the spontaneous symmetry breaking seen in the SM and a composite BEH boson.
II Implementing the Relative Time Symmetry